Dan tossed unbiased coin , on each toss he wins 50 pesos if head appears and losses 40 pesos if tail appears. What will be the expected value and variance of his gain?​


1
Expert's answer
2022-03-19T02:38:17-0400

Since the coin is unbiased, the probability of getting a head or a tail is both "=0.5"


The expected value is thus


"=(50\times 0.5)-(40\times0.5)"

"= 5" Pesos.


The variance


"\begin{matrix} x&P(x)&xP(x)&x^2P(x)\\-40&0.5&-20&800\\50&0.5&25&1250\\\sum &1&5&2050\end{matrix}"


"\sigma^2=\sum x^2P(x)-\sum (xP(x))^2"


"\sigma^2 =2050-25"

"\sigma^2=2025"